OpenStudy (anonymous):

Divide. Assume that no denominator equals zero. 8d^3-6c divided by c^3d^4/24 A. 64c^2/d B.32/c^4d C.16c^2/d D.c^2d^7/18 please help

Nnesha (nnesha):

\[\huge\rm \frac{ 8d^3 -6c }{ \frac{c^3d^4}{24}}\] like this ??

Nnesha (nnesha):

if top part is dividing by fraction so change it to multiplication multiply top part by the reciprocal of bottom part \[\huge\rm \frac{ \frac{a}{b} }{ \frac{x}{y} } \]\[\large\rm=\frac{ a }{ b } \times \frac{ y }{ x }\]